OpenStudy (anonymous):

A group of friends were voting for their favorite song: four-ninths chose song A, one-fifth of the remaining friends chose song B, and the rest chose song C. If there were eighteen friends in all, how many chose song C?

OpenStudy (anonymous):

@mr.mangle can you help

OpenStudy (anonymous):

A + B + C=18 A = (4/9)18 B = (1/5) (18-A) C= 18-A-B

OpenStudy (anonymous):

@texaschic101 can you help me

OpenStudy (texaschic101):

ok...there is 18 friends 4/9 chose song A 4/9(18) = 72/9 = 8 chose song A 1/5 of remaining friends chose song B so 1/5 of 18 - 8 = 1/5 of 10 = 1/5(10) = 10/5 = 2 chose song B the rest chose song C 18 - 8 - 2 = 10 - 2 = 8 chose song C any questions ?

OpenStudy (anonymous):

No Question But thank you so so much @texaschic101

OpenStudy (texaschic101):

no problem :)